# Hydrodynamics without Boost-Invariance from Kinetic Theory – From Perfect Fluids to Active Flocks

### v2

@@ I Introductory Remarks
The fact that one is forced to factor in momentum exchange with the reservoir if one is to arrive at a first-principles understanding of a generic boost non-invariant system – a flock of birds being the most relatable example – should be immediately apparent to anyone who has ever seen a flock change direction mid-air. By flapping their wings, momentum is transferred to the reservoir (air in this case), allowing birds to accelerate and change their direction of motion. Although obvious to the point that would justify it being taken as a given, such a starting vantage has yet to be formalized and developed in the context of flocking theory, particularly concerning active flocks where the individual birds (or boids) [^2] For the unfamiliar: Boid originates from the name of an artificial life program [27], where it’s a portmanteau of bird-oid for bird-like objects. It has also been noted to be the New Yorker pronunciation for ‘bird’ [28]. can themselves act as energy, momentum, and entropy sources and sinks, which can be implemented via collision terms consistent with these assumptions in any kinetic theory derivation. The treatment that follows represents initial steps in this direction.
@@ II From Kinetic Theory to Boost Non-Invariant Hydrodynamics
We begin our treatment by assuming the existence of a distribution function $f(t,{\mathbf{x}},\hat{{\mathbf{p}}})$, which is a function of time, space, and momentum. More precisely, $f$ counts the number of single-particle states. We assume that the system obeys classical statistics, that is, that quantum statistical effects can be neglected. Therefore, the particle number density is
$n(t,\mathbf{x})=\int\frac{d^{d}\hat{p}}{h^{d}}\,f(t,\mathbf{x},\hat{\mathbf{p}}).$ (6)
@@
where $\tau$ is some collision time (itself possibly space- and time-dependent). This form for the collision term also goes by the name of the single relaxation, or single collision time approximation. Plugging this into the right hand side of (11) indeed shows that when $f>f^{\text{eq}}$, the evolution equation of $f$ will tend to decrease $f$, and if $f<f^{\text{eq}}$, it will tend to increase it, tending towards $f^{\text{eq}}$ in both cases. For $f$ sufficiently close to $f^{\text{eq}}$ this is simply linear response in the space of distributions. Of course, one could consider a more complicated dependence on $f-f^{\text{eq}}$ so long as the collision functional satisfies three requirements [37]:
@@ II.2 First Moment: Pre-Euler Equation
We now make the crucial assumption that we can factorize the expectation value of $\hat{\rho}$ from any other expectation value. This assumption derives from rotational-invariance and the additional assumption that we can ignore sufficiently high velocity moments [^9] This relates to the manner in which we truncate the Vlasov hierarchy that would ordinarily go on indefinitely by assuming the ‘enslavement’ of higher order moments. We return to this point in Section II.3.. Rotational-invariance implies that $\hat{\rho}$ depends on the quantity $\hat{v}^{2}$, rather than on $\hat{{\mathbf{v}}}$ itself. Thus, if we can ignore third-order velocity moments, we can always factorize $\hat{\rho}$ out of any expectation value. Hence, equation (23) simplifies to
@@ II.3 Second Moment: the Generalized Euler Equation
where $P$ is some function of $v^{2}$, which we would like to interpret as pressure, but which, for now, we take to be some arbitrary function. This is the most general expression we can construct out of the expectation value $v^{i}$ that has the correct tensor structure and does not involve spatial gradients. Therefore,
@@ III Incorporating Inter-particle/boid Forces
We first emphasize that $\hat{\mathbf{\boldsymbol{\eta}}}$ is not to be thought of as a (global) mean field. Instead, à la Vicsek, it is to be thought of as a local average of the velocities of the nearby boids whose precise form remains to be specified. It is essentially a non-local quantity, but in a manner that is made transparent and tractable when written, without loss of generality, as the convolution of an influence kernel $K^{i}{}_{j}({\mathbf{r}},{\mathbf{r}}^{\prime})$ with the velocity field:
@@
We note that the influence kernel will have finite support for short-range forces, and its precise form serves as a functional parameterization of all possible two-body inter-boid or inter-particle forces at the microscopic, or kinetic theory level [^12] One can in principle include higher point interactions by adding additional quadratic and higher convolutions to (37) as desired, i.e.: $\hat{\eta}^{i}(\mathbf{r})=\int d^{d}r^{\prime}\,K^{i}{}_{j}(\mathbf{r},\mathbf{r}^{\prime})\,\hat{v}^{j}({\mathbf{r}}^{\prime})+\int d^{d}r^{\prime}\int d^{d}r^{\prime\prime}\,K^{i}{}_{jk}(\mathbf{r},\mathbf{r}^{\prime},\mathbf{r}^{\prime\prime})\,\hat{v}^{j}({\mathbf{r}}^{\prime})v^{k}({\mathbf{r}}^{\prime\prime})+...$ etc..
@@ IV The Hydrodynamics of Flocks
where we drop diffusion terms for the time being. We subscript the density and pressure referenced in (44) with zeroes to distinguish them from the corresponding quantities in the present treatment, since these are not the same: in [22], these quantities are taken to be independent of velocity, which, as discussed in previous sections, is not something one can take for granted on general kinematic grounds for a boost non-invariant fluid.
@@
In the context of particle physics effective field theories, the tendency for dimensionless parameters to either vanish or be close to unity in certain units (being exactly zero or unity when a symmetry is restored) is known as technical naturalness [44]. It is a statement about how hierarchies are preserved under renormalization group flow at long wavelengths, and boils down to the fact that these parameters are multiplicatively (rather than additively) renormalized as one flows to the infrared [^18] That is, the renormalization group equations for a given parameter is proportional to powers of the difference of that parameter from its symmetry restored value at any given renormalization group scale.. Whether the notion of technical naturalness transposes to the present context to yield the hierarchy (55) from a renormalization group analysis as suggested by our treatment in Appendix E is an important question we leave for the future. For the present, we note that under a minimal set of assumptions, a prediction of our treatment is that any active flock that satisfies the assumption of isotropy and inter-boid momentum conservation will satisfy the relations (55).
@@
The diffusion coefficients themselves are given by the parameter $\lambda$ that features in the force ansatz (77) and the parameters that define the functional form of the influence kernel, as illustrated in Table 1 for isotropic kernels (sampled from Table 2 in Appendix C). A non-zero $D_{B}$ diffusion coefficient can be generated through an anisotropic influence kernel (96), where furthermore, $D_{B}$ and $D_{T}$ relate to each other via the parameterized departure from isotropy.
@@ V Discussion and Outlook
The main insights that led us to our conclusions, from which our derivations depart ab initio from standard treatments (see e.g. [26][39]) involve:
@@
It is informative to compare our results to the process of course-graining a specific microscopic model with the standard Newtonian (i.e. Galilean boost invariant) kinematic relation between $\rho$ and $mn$, as performed in [39]. Here, one can derive the Dean equation from the Langevin equations describing the microscopic interactions, and systematically expand for the leading order modes at long wavelengths. The Toner-Tu theory which the Vicsek model flows to corresponds to a description where the following ratios are fixed as: $\lambda_{1}/\lambda_{2}=3/5$ and $\lambda_{2}/\lambda_{3}=-2$. Among other things, there is no scope for $\lambda_{2}$ to vanish, which is the condition assumed in [22] in order to derive relationships between scaling exponents. In our approach, $\lambda_{2}=0$ simply translates to $g$ being independent of $\nabla\cdot{\mathbf{v}}$ at leading order, albeit in a model which qualitatively differs from the Vicsek model in that it allows for longitudinal breathing modes via the parameterizations (36) and (37). Whether or not $\lambda_{2}=0$ also follows from a renormalization group analysis within our formalism applied to an influence kernel specific to the Vicsek model is an important question which behooves us to revisit.
@@ Appendix C The Influence Kernel and Local Mean-Field Velocity Matching
$\hat{{\mathbf{F}}}=-\lambda(\hat{{\mathbf{p}}}-\hat{{\mathbf{\boldsymbol{\eta}}}}),$ (77)
where $\lambda$ is a constant with dimensions of inverse time and ${\mathbf{\boldsymbol{\eta}}}$ is a “local mean-field momentum” term, which is a kind of average of the momenta of the nearby boids. With $\lambda>0$, this force tends to align and match an individual boid’s momentum with the local mean-field momentum of the nearby boids. We write the local mean-field momentum as the convolution of a spatially-averaging kernel $K^{i}{}_{j}({\mathbf{r}})$ with the momentum field:
@@
where, $\hat{{\mathbf{p}}}$ is technically not a function of ${\mathbf{r}}^{\prime}$, of course, since no expectation value has been taken yet. However, when we do take an expectation value over momentum, this factor will be a function of ${\mathbf{r}}^{\prime}$. The kernel measures the strength of the influence of one boid on another. Therefore, we take to calling $K^{i}{}_{j}({\mathbf{r}})$ the influence kernel.
@@
where $\widetilde{K}^{i}{}_{j}(\mathbf{k})$ and $\tilde{\hat{p}}^{j}(\mathbf{k})$ are the Fourier transforms of $K^{i}{}_{j}(\mathbf{r})$ and $\hat{p}^{j}(\mathbf{r}^{\prime})$, respectively. Finally, we may write this as
@@
    where $\Theta$ is the Heaviside theta function and $\kappa$ is some dimensionless constant setting the normalization of the influence kernel [^25] The form of this influence kernel is morally correct even though the force (77) is not quite the same as that of the Vicsek model. The latter assumes that the boids all have the same speed and that boid “$a$” aligns its orientation $\theta_{a}$ with boid “$b$” via an interaction of the form $\sin(\theta_{b}-\theta_{a})$. Thus, our starting ansatz for the force is not quite analogous to that of the Vicsek model. Nevertheless, the step-function form of the influence kernel, is correct..
    Dimensional analysis implies that every gradient factor must be multiplied by a length scale factor. For the kernel given in (81), for example, we know that $\widetilde{K}^{i}{}_{j}$ is not just a function of $-i\nabla$, but rather of $-ir_{0}\nabla$, for the same reason that $K^{i}{}_{j}$ is not just a function of ${\mathbf{r}}$, but of $\frac{{\mathbf{r}}}{r_{0}}$.
@@
    We may interpret the influence kernel as the Green’s function for some spatial differential operator ${\cal O}^{i}{}_{j}(\nabla)$:
@@
    where $\kappa$ is a constant with the same dimensions as ${\cal O}^{i}{}_{j}$. The operator ${\cal O}^{i}{}_{j}$ and the constant $\kappa$ may be rendered dimensionless by absorbing dimensions using a localization scale $r_{0}$, as done in (81). For instance, the example discussed in sec. III is the Green’s function for the (dimensionful) Helmholtz operator, ${\cal O}^{i}{}_{j}(-i\nabla)=\delta^{i}{}_{j}(\nabla^{2}+\mu^{2})$, which gives the famous Yukawa potential. As discussed in sec. III, $\kappa$ also sets the fluctuation scale that allows this formalism to capture stochastic effects, such as diffusion, without the need to introduce noise terms in the force. Indeed, at lowest non-trivial order in spatial gradients, the gradient expansion of $\widetilde{K}^{i}{}_{j}(-i\nabla)$ generates the two diffusion terms in the Toner-Tu equation: $\nabla^{2}v^{i}$ and $\nabla^{i}(\nabla\cdot{\mathbf{v}})$.
@@ C.1 Influence Kernels and Diffusion Coefficients in 2d
The Fourier transform of the influence kernel has a problem, however, since $k^{i}k_{j}$ can become arbitrarily negative, which makes the off-diagonal components of $\widetilde{K}^{i}{}_{j}$ unbounded. It also vanishes along the axes. There are many ways to remedy this, but all of them require breaking the nice tensorial structure of $\widetilde{K}^{i}{}_{j}$ to some extent. For example, we can simply assume that there is an off-diagonal $O(k^{4})$ term that solves this problem, such as $-\epsilon(r_{0}^{2}k^{2})^{2}$. This breaks the nice tensorial structure by a small amount if $\epsilon\ll 1$ in exchange for bounding the off-diagonal components of $\widetilde{K}^{i}{}_{j}$. This has the unfortunate consequence that we cannot actually write down the off-diagonal components of the influence kernel in space: without the $\epsilon$ term, the inverse Fourier transform is not well-defined, and with the $\epsilon$ term, the inverse Fourier transform does not have a closed analytic form. Nevertheless, we can write down the diagonal components of the influence kernel in space:
$K^{x}{}_{x}({\mathbf{r}})$ $=\kappa\,\frac{e^{-\frac{1}{2(1+2\chi)}(\frac{x}{r_{0}})^{2}}}{\sqrt{2\pi(1+2\chi)}\,r_{0}}\cdot\frac{e^{-\frac{1}{2}(\frac{y}{r_{0}})^{2}}}{\sqrt{2\pi}\,r_{0}},$ (97a)
$K^{y}{}_{y}({\mathbf{r}})$ $=\kappa\,\frac{e^{-\frac{1}{2}(\frac{x}{r_{0}})^{2}}}{\sqrt{2\pi}\,r_{0}}\cdot\frac{e^{-\frac{1}{2(1+2\chi)}(\frac{y}{r_{0}})^{2}}}{\sqrt{2\pi(1+2\chi)}\,r_{0}}.$ (97b)
@@
$\hat{{\mathbf{F}}}=-\frac{\hat{{\mathbf{p}}}}{\tau^{\prime}}-\lambda(\kappa\hat{{\mathbf{p}}}-\hat{{\mathbf{\boldsymbol{\eta}}}}),$ (98)

### v3

@@ I Introductory Remarks
The fact that one is forced to factor in momentum exchange with the reservoir if one is to arrive at a first-principles understanding of a generic boost non-invariant system – a flock of birds being the most relatable example – should be immediately apparent to anyone who has ever seen a flock change direction mid-air. By flapping their wings, momentum is transferred to the reservoir (air in this case), allowing birds to accelerate and change their direction of motion. Although obvious to the point that would justify it being taken as a given, such a starting vantage has yet to be formalized and developed in the context of flocking theory, particularly concerning active flocks where the individual birds (or boids) can themselves act as energy, momentum, and entropy sources and sinks, which can be implemented via collision terms consistent with these assumptions in any kinetic theory derivation. For the unfamiliar: the term boid originates from the name of an artificial life program [27], where it’s a portmanteau of bird-oid for bird-like objects. It has also been noted to be the New Yorker pronunciation for ‘bird’ [28].
@@ II From Kinetic Theory to Boost Non-Invariant Hydrodynamics
[…]
@@
where $\tau$ is some collision time (itself possibly space- and time-dependent). [^7] The term “collision time” here refers to the time between collisions, as opposed to the duration of a collision. The time $\tau$ is sometimes also referred to as a “relaxation time.” As usual, we assume that the duration of a collision can be ignored: that time scale is much shorter than $\tau$, the time between collisions. This assumption implicitly involves both density and interaction strength: the higher both of those are, the lower we expect $\tau$ to be, and the more one has to be careful about the approximation. This form for the collision term also goes by the name of the single relaxation, or single collision time approximation. Plugging this into the right hand side of (11) indeed shows that when $f>f^{\text{eq}}$, the evolution equation of $f$ will tend to decrease $f$, and if $f<f^{\text{eq}}$, it will tend to increase it, tending towards $f^{\text{eq}}$ in both cases. For $f$ sufficiently close to $f^{\text{eq}}$ this is simply linear response in the space of distributions. Of course, one could consider a more complicated dependence on $f-f^{\text{eq}}$ so long as the collision functional satisfies three requirements [37]:
@@ II.2 First Moment: Pre-Euler Equation
We now make the crucial assumption that we can factorize the expectation value of $\hat{\rho}$ from any other expectation value. This assumption derives from rotational-invariance and the additional assumption that we can ignore sufficiently high velocity moments [^9] This relates to the manner in which we truncate the Vlasov hierarchy that would ordinarily go on indefinitely by assuming the ‘enslavement’ of third- and higher-order moments. Generally, one assumes the enslavement of all moments of higher order than the one of interest. We are interested in the two-point function, and thus we will truncate at third order. This is the usual procedure (see [37]). We return to this point in Section II.3.. Rotational-invariance implies that $\hat{\rho}$ depends on the quantity $\hat{v}^{2}$, rather than on $\hat{{\mathbf{v}}}$ itself. Thus, if we can ignore third-order velocity moments, we can always factorize $\hat{\rho}$ out of any expectation value. Hence, equation (23) simplifies to
@@ II.3 Second Moment: the Generalized Euler Equation
where $P$ is some function of $v^{2}$, which we would like to interpret as pressure, but which, for now, we take to be some arbitrary function. This is a standard expression used in the literature; it is the most general expression we can construct out of the expectation value $v^{i}$ that has the correct tensor structure and does not involve spatial gradients. The expression is understood to hold in the bulk of the fluid, where equilibrium conditions justify excluding spatial gradients. Of course, in the vicinity of boundaries spatial gradients may naturally arise and the situation will be more complicated. Therefore,
@@ III Incorporating Inter-particle/boid Forces
Furthermore, we note that unless we impose restrictions on the form of the distribution function, the parameterization (36) as expressed departs from anything that could reproduce the Vicsek model which considers normalized velocities, and consequently only aligns directions to a local averaged field (with window function-like support). The ansatz (36) instead allows for longitudinal breathing modes corresponding to boids or particles getting ahead of or catching up with the local flow. In order to reproduce anything that could resemble the Vicsek model, one would either have to restrict the distribution function to have no support outside a momentum shell that corresponds to the constraint $|{\mathbf{v}}|=1$ (or some other constant), rendered non-trivial because of the relation (3). The net result would mimic adding additional contact force terms to (36) to enforce the momentum shell constraints. Equivalently, one can rederive the Vlasov hierarchy from the beginning but replacing the Poisson brackets implicit in the Boltzmann equation [^11] Where we note that in the case that inter-boid forces derive from an underlying symplectic structure, (II) can be rewritten as $\frac{df}{dt}=\frac{\partial f}{\partial t}+\frac{dx^{i}}{dt}\frac{\partial f}{\partial x^{i}}+\frac{d\hat{p}_{i}}{dt}\frac{\partial f}{\partial\hat{p}_{i}}\equiv\frac{\partial f}{\partial t}+\{\mathcal{H},f\}$ where $\mathcal{H}$ is the generator of time evolution, and where $\{~,~\}$ denotes the corresponding Poisson bracket. (II) with the corresponding Dirac brackets, and considering the equivalent constrained Hamiltonian evolution [41, 42, 43]. We leave this as an exercise for the future, opting for the simple unconstrained parameterization (36) in what follows, as longitudinal breathing modes in a flock are evidently well motivated from physical considerations, and moreover, is straightforward to deal with in the calculations to follow. We could, of course, consider more general possibilities than the ansatz (36) as briefly discussed in Appendix C, however, this simple parameterization already possesses a rich phenomenology which warrants an expanded discussion. Of particular importance, is how microscopic stochasticity manifests at long wavelengths even as the addition of noise terms to (36) drops out of hydrodynamic averages.
@@
We note that the influence kernel will have finite support for short-range forces, and its precise form serves as a functional parameterization of all possible two-body inter-boid or inter-particle forces at the microscopic, or kinetic theory level [^12] One can in principle include higher point interactions by adding additional quadratic and higher convolutions to (37) as desired, i.e.: $\hat{\eta}^{i}(\mathbf{r})=\int d^{d}r^{\prime}\,K^{i}{}_{j}(\mathbf{r},\mathbf{r}^{\prime})\,\hat{v}^{j}({\mathbf{r}}^{\prime})+\int d^{d}r^{\prime}\int d^{d}r^{\prime\prime}\,K^{i}{}_{jk}(\mathbf{r},\mathbf{r}^{\prime},\mathbf{r}^{\prime\prime})\,\hat{v}^{j}({\mathbf{r}}^{\prime})v^{k}({\mathbf{r}}^{\prime\prime})+$... etc.. The influence kernel captures only the inter-boid interactions, not the interactions between the boids and the environment. As stated earlier, we assume translation-invariance of the inter-boid interactions, which is why $K$ is a function only of $|{\mathbf{r}}-{\mathbf{r}}^{\prime}|$. In addition, we will mostly assume that the inter-boid interactions are also isotropic, which is why $K^{i}{}_{j}\propto\delta^{i}{}_{j}$. Nevertheless, we show in Appendix C how anisotropic terms in the influence kernel can naturally lead to diffusion terms in the Euler equation.
@@ IV The Hydrodynamics of Flocks
where we drop diffusion terms for the time being. [^15] Implicit in this equation is the assumption of the analyticity of the small-$\mathbf{v}$} expansion as, otherwise, scalar quantities in this equation could have depended on $v=|{\mathbf{v}}|$ instead of $v^{2}=|{\mathbf{v}}|^{2}$. We assume this throughout as well. We subscript the density and pressure referenced in (44) with zeroes to distinguish them from the corresponding quantities in the present treatment, since these are not the same: in [22], these quantities are taken to be independent of velocity, which, as discussed in previous sections, is not something one can take for granted on general kinematic grounds for a boost non-invariant fluid.
@@
In the context of particle physics effective field theories, the tendency for dimensionless parameters to either vanish or be close to unity in certain units (being exactly zero or unity when a symmetry is restored) is known as technical naturalness [45]. It is a statement about how hierarchies are preserved under renormalization group flow at long wavelengths, and boils down to the fact that these parameters are multiplicatively (rather than additively) renormalized as one flows to the infrared [^19] That is, the renormalization group equations for a given parameter is proportional to powers of the difference of that parameter from its symmetry restored value at any given renormalization group scale.. Whether the notion of technical naturalness transposes to the present context to yield the hierarchy (55) from a renormalization group analysis as suggested by our treatment in Appendix E is an important question we leave for the future. In this regard, we should note that the question of a proper and complete renormalization group analysis for the theory of flocks is not settled (see, for example, [46, 47, 48, 49]). For the present, we note that under a minimal set of assumptions, a prediction of our treatment is that any active flock that satisfies the assumption of isotropy and inter-boid momentum conservation will satisfy the relations (55).
@@
The diffusion coefficients themselves are given by the parameter $\lambda$ that features in the force ansatz (77) and the parameters that define the functional form of the influence kernel, as illustrated in Table 1 for isotropic kernels (sampled from Table 2 in Appendix C). A non-zero $D_{B}$ diffusion coefficient can be generated through an anisotropic influence kernel (96), where furthermore, $D_{B}$ and $D_{T}$ relate to each other via the parameterized departure from isotropy. Finally, noise can be added in by hand, as is done in Toner-Tu theory.
@@ V Discussion and Outlook
The main insights that led us to our conclusions, from which our derivations depart ab initio from standard treatments (see e.g. [26, 39]) involve:
@@
It is informative to compare our results to the process of course-graining a specific microscopic model with the standard Newtonian (i.e. Galilean boost invariant) kinematic relation between $\rho$ and $mn$, as performed in [39]. Here, one can derive the Dean equation from the Langevin equations describing the microscopic interactions, and systematically expand for the leading order modes at long wavelengths. The Toner-Tu theory which the Vicsek model flows to corresponds to a description where the following ratios are fixed as: $\lambda_{1}/\lambda_{2}=3/5$ and $\lambda_{2}/\lambda_{3}=-2$. In [40], one can also find a derivation of Toner-Tu theory from a coarse-graining of the Vicsek model via a different approach, and it would be useful to compare the two methods to each other. However, the comparison would be tangential to the present approach as both start from a specific model rather than a generic model parameterization, and differ in assuming the usual Newtonian relation between momentum density, number density, particle mass, and velocity. In both of the aforementioned approaches, among other things, there is no scope for $\lambda_{2}$ to vanish, which is the condition assumed in [22] in order to derive relationships between scaling exponents. In our approach, on the other hand, $\lambda_{2}=0$ simply translates to $g$ being independent of $\nabla\cdot{\mathbf{v}}$ at leading order, albeit in a model which qualitatively differs from the Vicsek model in that it allows for longitudinal breathing modes via the parameterizations (36) and (37). Whether or not $\lambda_{2}=0$ also follows from a renormalization group analysis within our formalism applied to an influence kernel specific to the Vicsek model is an important question which behooves us to revisit.
@@ Appendix C The Influence Kernel and Local Mean-Field Velocity Matching
[…]
@@
[…]
@@
[…]
@@
    where $\Theta$ is the Heaviside theta function and $\kappa$ is some dimensionless constant setting the normalization of the influence kernel [^26] The form of this influence kernel is morally correct even though the force (77) is not quite the same as that of the Vicsek model. The latter assumes that the boids all have the same speed and that boid “aa” aligns its orientation $\theta_{a}$ with boid “bb” via an interaction of the form $\sin(\theta_{b}-\theta_{a})$. Thus, our starting ansatz for the force is not quite analogous to that of the Vicsek model. Nevertheless, the step-function form of the influence kernel, is correct..
    Dimensional analysis implies that every gradient factor must be multiplied by a length scale factor. For the kernel given in (81), for example, we know that $\widetilde{K}^{i}{}_{j}$ is not just a function of $i\nabla$, but rather of $ir_{0}\nabla$, for the same reason that $K^{i}{}_{j}$ is not just a function of $\mathbf{r}$}, but of $\frac{{\mathbf{r}}}{r_{0}}$.
@@
[…]
@@
[…]
@@ C.1 Influence Kernels and Diffusion Coefficients in 2d
The Fourier transform of the influence kernel has a problem, however, since $k^{i}k_{j}$ can become arbitrarily negative, which makes the off-diagonal components of $\widetilde{K}^{i}{}_{j}$ unbounded. It also vanishes along the axes. There are many ways to remedy this. We will discuss two methods below.
Method 1: Add an $O(k^{4})$ term, such as $\epsilon(r_{0}^{2}k^{2})^{2}$, to the off-diagonal components of $\widetilde{K}^{i}{}_{j}$. This breaks the nice tensorial structure by a small amount if $\epsilon\ll 1$ in exchange for bounding the off-diagonal components of $\widetilde{K}^{i}{}_{j}$. This has the unfortunate consequence that we cannot actually write down the off-diagonal components of the influence kernel in space: without the $\epsilon$ term, the inverse Fourier transform is not well-defined, and with the $\epsilon$ term, the inverse Fourier transform does not have a closed analytic form. Nevertheless, we can write down the diagonal components of the influence kernel in space:
[…]
@@
Method 2: A simpler method would be to write $\widetilde{K}^{i}{}_{j}$ as
$\widetilde{K}^{i}{}_{j}({\mathbf{k}})=\kappa\bigl{(}\delta^{i}_{j}-\chi r_{0}^{2}k^{i}k_{j}\bigr{)}e^{-\frac{1}{2}(r_{0}k)^{2}},$ (98)
instead of (93). In real space, this is a superposition of an isotropic $\delta^{i}_{j}$ term as well as an anisotropic Hessian-type term [^27] Thanks to the referee for pointing out the possibility of a Hessian term in the influence kernel.:
$K^{i}{}_{j}({\mathbf{r}})=\frac{\kappa}{2\pi r_{0}^{2}}\biggl{[}\delta^{i}_{j}+\chi r_{0}^{2}\frac{\partial}{\partial r_{i}}\frac{\partial}{\partial r^{j}}\biggr{]}e^{-\frac{1}{2}(\frac{r}{r_{0}})^{2}},$ (99)
which evaluates to
$K^{i}{}_{j}({\mathbf{r}})=\frac{\kappa}{2\pi r_{0}^{2}}\biggl{[}(1-\chi)\delta^{i}_{j}+\chi\frac{r^{i}r_{j}}{r_{0}^{2}}\biggr{]}e^{-\frac{1}{2}(\frac{r}{r_{0}})^{2}}.$ (100)
@@
[…]